# Emergent Coordination in Multi-Agent Language Models

### v1

@@ 3.2 Main Experiments
For the main set of experiments, we chose groups of $N=10$ because that appears to be the most difficult setup, and a temperature of 1, and used OpenAI’s gpt-4.1-2025-04-14 (OpenAI, 2025) and Meta’s Llama-3.1-8B (Meta, 2024) with the prompts shown in Appendix A.1. We query gpt-4.1 through the OpenAI API and Llama through locally hosted Ollama with a NVIDIA RTX A6000. We replicate each group experiment 200 times per treatment condition (600 experiments total). Overall success rate is not significantly different across the interventions (Figure 2a).
@@
Plain: The control condition uses only instructions for the guessing game, similar to those used in human subject experiments (Goldstone et al., 2024). Persona: For the persona condition, we follow recent insights into the use of personas within the LLM community (Chen et al., 2024) and ensure personas have relevant attributes: name, gender, age, occupation or skill background, pronouns, personality traits (Goldberg, 1992), and personal values (Cieciuch & Davidov, 2012). ToM: The ToM condition instructs agents to think about what other agents might do, and how their actions might affect the group outcome, inspired by chain-of-thought reasoning (Wei et al., 2022). See Appendix A.1 for all prompts, full persona generation recipes, and example persona.
@@ 4.1 Emergent Synergy
We repeat analyses using the emergence capacity criterion. We find it is significantly above its null across conditions; results hold under residualization and alternative entropy corrections (Figure 2c; see Appendix for details). Taken together, across two different tests (one using a macro signal the other using future system behavior) and across a variety of robustness tests we find evidence of dynamic emergence capacity in multi-agent LLM systems (answering RQ1).
Next, we explore whether agents evolve specialized roles & identities. Are there detectable between-agent differences in either the level of their contribution to the group guess or their temporal evolution (i.e., learning rate)? Using the hierarchical model comparison test we find that agents in may groups have differentiated identities (Figure 3a; a Fisher test for joint $p$-value evidence is again highly significant). There is substantially more differentiation among agents in the Persona condition, and even more in the ToM condition. Inspecting the reasoning traces of agents, they often contain references to the “personal experience” of their assigned persona such as “In my experience (whether it’ corraling cattle or wrangling numbers), starting toward the middle gives the group the best shot to zero in after the first feedback” (see Appendix for additional sample quotes).
4.2 What does Emergence Allow?
Internal Organization of Groups beyond Synergy.
What functional differences—if any—does emergence allow? To explore if the internal organization of behavior in groups is different across treatment conditions, we contrast total mutual information ($I_{3}$) with information contained in triplets ($G_{3}$). We find significant above zero dynamic emergent synergy in triplets that is not reducible to pairs in all three treatment conditions (again the $p$-value test shows 20% of groups below 0.05 and a highly significant Fisher test). This provides further evidence for multi-agent LLM system capacity for emergence. $G_{3}$ is significantly lower in the ToM condition compared to the Plain condition (two-sample KS test $p=0.036$) but not different from the Persona condition ($p=0.842$; Figure 3b). However, we find significantly higher $I_{3}$ in the ToM condition compared to both other conditions (both $p<0.001$; Figure 3c). Finally, we test how many groups have significantly positive $I_{3}$. We find ToM condition has substantially more such groups (Figure 3d). In Appendix A.9 we provide additional robustness tests ruling out heterogeneous learning rates as the sole source of differentiation and provide additional evidence for stable agent identities that persist over time.
Increased Performance.
We use regression analysis to explore the joint effect of synergy and redundancy (based on emergence capacity formulation) while carefully controlling for endogeneity of variables measuring emergent properties and performance and controlling for sample selection bias with stabilized inverse probability weighting (see Appendix A.10 for method details). On their own, higher levels of either synergy or redundancy do not predict success. However, when both are present, performance improves significantly (significant interaction with $\beta=0.24$; $p=0.014$). In marginal-effect terms, redundancy amplifies the benefit of synergy on the log-odds scale by 27%; and vice versa synergy amplifies benefits of redundancy by 27%. This pattern implies that systems benefit when redundant pathways create goal alignment, while synergistic interactions extract novel, non-overlapping information—together enabling higher overall performance. We complement the regression analysis with causal mediation analysis (Imai et al., 2010). While reaching only marginal significant levels the effect is consistent: the ToM treatment causally increases performance indirectly by increasing synergy (ACME =0.034=0.034 [95%CI: 0.0000.07-0.000-0.07], $p=0.053$). This aligns with the interpretation that performance benefits emerge when systems achieve both redundancy (aligned toward a common goal) and synergistic integration (differentiated, complementary roles)—a form of functional complexity (Varley et al., 2023; Luppi et al., 2024).
@@ 4.2 What does Emergence Allow?
Figure 3: a) Agent differentiation using hierarchical mixed model comparison (counting groups in which at least one test (different intercepts or slopes) is below $p<0.05$. b) Total time-delayed mutual information of triplets ($I_{3}$). c) Dynamic emergent synergy ($G_{3}$). d) ToM-prompt condition has substantially more groups with significant $I_{3}$ content (above 0). In panel a) and d), error bars show Wilson confidence intervals for binary data. Stars indicate significance level of test for equal proportion. Panel b) and c) shows raw data with Jeffreys’ correction. Data are Winsorized at the 1st and 99th percentiles for visual clarity. Stars indicate significance level of Wilcoxon test. Notes: *** $p<0.001$; ** $p<0.01$; * $p<0.05$.
Together, the results suggest a nuanced picture of emergence in multi-agent systems. Persona and ToM produce agents with distinct, stable identities—moving agents on the differentiated-undifferentiated axis in Figure 1a. This differentiation shapes the space in which other agents can adapt, and specialization by one agent constrains the (useful) actions of the others (agents mutually constrain each other; cf. Tollefsen et al., 2013)—shifts on the interdependent-complementary axis. When combined with ToM, mutual adaptation both strengthens agent differentiation (creating the foundation for possible synergistic complementarity) while also increasing mutual alignment on shared goals (the target signal)—shifts on the integration axis. Differentiation enables synergy, but without redundancy and integration, synergy alone does not translate into better collective performance. Agents in the ToM condition develop specialized, complementary roles while achieving stronger shared goal alignment: they are simultaneously more differentiated yet also more integrated along highly overlapping task-relevant information which the capacity for synergy can now exploit in productive ways (answering RQ2 and RQ3).
4.3 Experiments using Llama-3.1-8B
Repeating the experiments using another class of LLM with 8B parameters (Llama-3.1) we find multi-agent systems are generally not able to solve the task (at least not with the identical prompt as used in the other experiments). Only 10% of groups manage to solve the task (Plain: 10.5%; Persona: 14%; ToM: 5.5%). Model capacity to support reasoning about other agents seems crucial and despite the instruction to do so, Llama-3.1-8B agents fail executing it effectively. Groups in the ToM condition perform significantly worse than in the Plain condition (test for equality of proportions $p=0.007$) thus reversing the findings of the more capable gpt-4.1 agents. Results regarding emergence capacity are mixed (see Appendix).
The strong oscillating pattern in guesses materializes as high predictive power over the group outcome (15.8% groups show significant $I_{3}$ compared to only 2% in gpt-4.1 groups) but much lower $G_{3}$ synergy between agents (significant synergy present in 2.6% Llama groups compared with 20% of gpt-4.1 groups). There is extremely strong coupling across time (current individual states predict future joint states well as agents oscillate) but very little cross-agent complementarity. This suggests that while multi-agent systems with lower-capacity LLMs still show some signs of emergence, it is mostly spurious across the temporal dimension rather than the productive cross-agent dimension found in gpt-4.1 agents. Lower ToM reasoning reasoning capacity in Llama-3.1-8B (Xiao et al., 2025) may explain the lack of useful, goal-directed cross-agent synergy. Capacity-matched prompt tuning (e.g., lower temperature or simplified instructions) might yield different results.
@@ 5 Related Work
Shortly after the release of ChatGPT, the community began exploring multi-agent LLM systems, where multiple LLM agents interact with each other. One early example is Park et al. (2023), who simulated a small population of “individuals” in a The Sims-style environment with friends, houses, and jobs, showing the emergence of rich social behaviors such as coordinating a Valentine’s Day party. This was followed by multi-agent systems for complex tasks such as software development (Chen et al., 2023; Hong et al., 2024; Qian et al., 2023), healthcare (Li et al., 2024a), and other domains (Subramaniam et al., 2025). Federated systems or “society of models” also gained attention (Juneja et al., 2024; Li et al., 2024b). These systems often achieve significant performance gains over single-agent baselines (Wu et al., 2023; Chen et al., 2023; Li et al., 2024b; Tao et al., 2024).
A recurring explanation for such gains invokes “greater-than-the-sum-of-its-parts” effects (Chen et al., 2023, p.1), often attributed to division of labor among differentiated agents (Subramaniam et al., 2025). Yet while this work is “inspired by human group dynamics” (Chen et al., 2023, p.1), it rarely evaluates whether the same principles that underlie effective (and ineffective) human groups also emerge in multi-agent systems. Two gaps stand out. First, human groups often coordinate by developing role specialization (Goldstone et al., 2024). Multi-agent systems often intuitively induce role specialization (such as “programmer”, “tester”, and “CEO”; Qian et al., 2023) yet have not systematically evaluated the impact on emergent shared cognition beyond win-rate comparisons. Second, groups only outperform individuals if their members specialize in contributing different pieces of information to the group task (Theiner, 2013; DeChurch & Mesmer-Magnus, 2010; Fazelpour & De-Arteaga, 2022). Effective groups balance complementarity—members contributing distinct information—with redundancy, the alignment on shared goals (Theiner, 2013; DeChurch & Mesmer-Magnus, 2010; Fazelpour & De-Arteaga, 2022; Luppi et al., 2024; Tollefsen et al., 2013). Too much of either undermines performance. Unlike win-rate–centric evaluations, we propose falsification tests and null modeling, thus extending multi-agent frameworks with formal theory. While these principles likely apply to LLM-based collectives given the universality of the integration-segregation tradeoff (Tononi et al., 1994), it remains unclear whether agents develop differentiated roles, whether such roles complement each other, how to steer it with prompts, and what role ToM capacity of models plays for collaboration (Westby & Riedl, 2023; Kleiman-Weiner et al., 2025; Riedl & Weidmann, 2025).
@@ 6 Conclusion
Collectives of LLM agents are emerging as a powerful new paradigm, capable of tackling tasks that exceed the reach of any single model. Yet raw capability alone may not determine their effectiveness. Effective multi-agent systems depend on acting as integrated, cohesive units as opposed to loose aggregations of individual agents. In collective intelligence, this distinction often separates groups that merely average their parts from those that achieve true synergy (Riedl et al., 2021). This paper asks whether such coordination dynamics also matter for LLM collectives, what functional benefits they have, and whether we can design prompts, roles, or reasoning structures that foster such integrated yet synergistic behavior. The framework we developed connects human group cognition theory to LLM multi-agent systems, offering a novel conceptual bridge. Methodologically, we demonstrate how to operationalize group-level synergy in AI collectives using quantitative information-theoretic measures. We also provide specific design principles for creating productive LLM collectives that are aligned on “shared goals,” which can inform multi-agent orchestration tools, cooperative AI, and mixed human–AI teamwork. This work extends ToM research beyond standard false-belief tests to realistic collaborative tasks (Shapira et al., 2024; 2023). Those insights will help us understand when multiple LLMs working together will be beneficial, why that is, and inform their design.
We note that evidence of higher-order synergy should not be interpreted as implying sophisticated cognition or consciousness. As conceptualized here, synergy is a structural property of part-whole relationships within multi-agent interactions. Such higher-order structures are known to arise in simple systems, including agents governed by reinforcement learning (Fulker et al., 2024) or via basic nonlinearities in contagion processes (Iacopini et al., 2019; Lee et al., 2025). Without attributing human-like cognition to the agents, the patterns of interaction we observe mirror well-established principles of collective intelligence in human groups (Riedl et al., 2021; DeChurch & Mesmer-Magnus, 2010): effective multi-agent performance requires both alignment on shared objectives and complementary contributions across members.
@@
This work focuses on developing a framework, characterizing emergence, and localizing it. While we explore the degree of goal alignment via a performance-related macro signal and performance effects of “early synergy”, the analysis linking synergy and redundancy to performance is challenging because synergy and redundancy are often co-dependent and emerge alongside performance. Future work should aim to more directly connect measures of synergy with performance. Despite a significant number of robustness tests, estimating entropy and establishing synergy is difficult. Furthermore, this work established results only using a single task—albeit one that is uniquely suited to studying synergy given how closely it mirrors ideal case of complementary behavior in a minimalist setting. More work will be necessary to convincingly establish when and how multi-agent LLM systems develop emergency synergy across tasks and measures. Given the small-data setting some of our information theoretic measures are limited to computing dynamic emergence on order $k=2$ which is bound to miss synergy of a higher order. Finally, this work points to the importance of ToM. LLMs capacity to act in a ToM-like manner appears crucial to achieve functional alignment in multi-agent systems. This should provide additional motivation for research on ToM in LLMs.

### v2

@@ 3.2 Main Experiments
For the main set of experiments, we chose groups of $N=10$ because that appears to be the most difficult setup and a temperature of $T=1$. We again used OpenAI’s GPT-4.1 (version date 2025-04-14; OpenAI, 2025) with the prompts shown in Appendix A.1. We replicate each group experiment 200 times per treatment condition (600 experiments total). Overall success rate is not significantly different across the interventions (Figure 2a). We perform robustness tests using four other models: Meta’s Llama-3.1-8B and Llama-3.1-70B-Instruct (Meta, 2024), Google’s Gemini 2.0 flash (DeepMind, 2025), and Alibaba’s Qwen3 235B A22B Instruct 2507 (Team, 2025). We query GPT-4.1 through the OpenAI API, Qwen throgh Cerebras, Gemini through OpenRouter, and Llama through locally a hosted vLLM on a NVIDIA RTX Pro 6000 Blackwell.
@@
Plain: The control condition uses only instructions for the guessing game, similar to those used in human subject experiments (Goldstone et al., 2024). Persona: For the persona condition, we follow recent insights into the use of personas within the LLM community (Chen et al., 2024) and ensure personas have relevant attributes: name, gender, age, occupation or skill background, pronouns, personality traits (Goldberg, 1992), and personal values (Cieciuch and Davidov, 2012). ToM: The ToM condition instructs agents to think about what other agents might do, and how their actions might affect the group outcome. This is akin to introducing social skills as defined in the human capital literature (Deming, 2022). See Appendix A.1 for all prompts, full persona generation recipes, and example persona.
@@ 4.1 Emergent Synergy
We repeat analyses using the emergence capacity criterion. We find it is significantly above its null across conditions (Figure 2c); results hold under residualization and alternative entropy corrections (see Appendix for details). Taken together, across two different tests (one using a macro signal the other using future system behavior) and across a variety of robustness tests we find evidence of dynamic emergence capacity in multi-agent LLM systems (answering RQ1).
4.2 Dynamical Mechanisms
How is this emergence dynamically maintained and where is it located within the groups? We find time-trend demeaned bias-corrected $I_{3}$ around 0 in the Plain (Wilcoxon $p$-value 0.974) and Persona condition ($p=0.846$) indicating groups that fail to escape chaotic, uncoordinated states. In contrast, the ToM shows significant positive mutual information ($p=3.5\times 10^{-14}$), collapsing the group’s behavior into a predictable macro-state (Figure A2). The number of groups with significantly positive $I_{3}$ is substantially higher in the ToM condition (Figure 3a), consistent with enhanced collective intelligence. To formalize this transition dynamic, we compute the Total Stability of the system (computed as $I_{3}$ normalized by the entropy of the macro-signal) which serves as a proxy for the Lyapunov stability of the collective state (Haddad and Chellaboina, 2008). We find Total Stability indistinguishable form zero in the Plain (Wilcoxon $p$-value $0.976$) and Persona ($p$-value = 0.858) indicating a “gaseous” state where groups fail to settle into a stable attractor. However, the ToM intervention induces a sharp increase in Total Stability ($p$-value = $2.9\times 10^{-14}$). Theoretically, this suggests the ToM prompt acts as a control parameter, steering the multi-agent system from a chaotic regime into a deep basin of attraction where collective behavior stabilizes and information processing capacity spikes: turning the system from a collection of individuals into a collective group.
Decomposition this stability reveals the specific mechanism of coordination. While the system exhibits emergence capacity and practical emergence, we find the stability in group behavior is supported by pairwise alignment rather than irreducible triplet complexity. Specifically, triadic information gain ($G_{3}$) is around 0 in Persona and ToM (Figure 3c). The Plain conditions shows small positive $G_{3}>0$ (Wilcoxon $p$-value = $0.026$) but given the near-zero Total Stability, this likely reflects transient stochastic correlations and oscillation rather than sustained coordination. In dynamical systems terms, the ToM prompt creates a deep basin of attraction where agents converge on a shared schematic (or Schelling point), making the collective state robust but informationally redundant. This mirrors “Mean Field” coupling in physics: because agents receive only global feedback and cannot observe individual peers, they couple to the aggregate signal rather than forming distinct local bonds. In such an environment, attempting complex higher-order synergy would be fragile; instead, the system converges on the efficient solution of dense pairwise alignment to the Mean Field.
@@ 4.2 Dynamical Mechanisms
Figure 3: a) ToM-prompt condition has substantially more groups with significant $I_{3}$ content (above 0). b) Total Stability (time-delayed mutual information of triplets ($I_{3}$ normalized by macro-signal entropy, bias corrected). c) Information gain of triads over most informative dyad ($G_{3}$, bias corrected). d) Agent differentiation using hierarchical mixed model comparison (counting groups in which at least one test (different intercepts or slopes) is below $p<0.05$. In panel a) and d), error bars show Wilson confidence intervals for binary data. Stars indicate significance level of test for equal proportion. Panel b) and c) shows bias corrected data with Jeffreys’ prior. Data are Winsorized at the 1st and 99th percentiles for visual clarity. Stars indicate significance level of Wilcoxon test. Notes: *** $p<0.001$; ** $p<0.01$; * $p<0.05$.
Next, we explore whether agents evolve specialized roles and identities. Are there detectable between-agent differences in either the level of their contribution to the group guess or their temporal evolution (i.e., learning rate)? Using the hierarchical model comparison test we find that agents in may groups have differentiated identities (Figure 3a; a Fisher test for joint $p$-value evidence is again highly significant). There is substantially more differentiation among agents in the Persona condition, and even more in the ToM condition. Inspecting the model output of agents, they often contain references to the “personal experience” of their assigned persona such as “In my experience (whether it’s corraling cattle or wrangling numbers), starting toward the middle gives the group the best shot to zero in after the first feedback” (see Appendix for additional sample quotes). In Appendix A.10 we provide additional robustness tests ruling out heterogeneous learning rates as the sole source of differentiation and provide additional evidence for stable agent identities that persist over time. That is, in the Plain condition, differentiation results only from idiosyncratic noise and transient drift introduced by probabilistic LLM responses. With Persona, agents gain stable, identity linked behavioral preferences that refine their differentiation. The addition, ToM sharpens role differentiation through feedback loops from mutual predictive modeling by conditioning on the public history (which functions as a common ground and coordination device; Lewis, 1969; Skyrms, 2010). This converts small persona-induced asymmetries into stable, self-reinforcing roles, creating the basis for strategic complementarity and stable alignment (answering RQ2 and RQ3)
4.3 Functional Role of Emergence in Group Performance
What functional differences—if any—does emergence allow? We use regression analysis to explore the joint effect of synergy and redundancy (based on emergence capacity formulation) while carefully controlling for endogeneity of variables measuring emergent properties and performance and controlling for sample selection bias with stabilized inverse probability weighting (see Appendix A.11 for method details). On their own, higher levels of either synergy or redundancy do not predict success. However, when both are present, performance improves significantly (significant interaction with $\beta=0.24$; $p=0.014$). In marginal-effect terms, redundancy amplifies the benefit of synergy on the log-odds scale by 27%; and vice versa synergy amplifies benefits of redundancy by 27%. This pattern implies that systems benefit when redundant pathways create goal alignment, while synergistic interactions extract novel, non-overlapping information—together enabling higher overall performance. We complement the regression analysis with causal mediation analysis (Imai et al., 2010). While reaching only marginal significant levels the effect is consistent: the ToM treatment causally increases performance indirectly by increasing synergy (ACME =0.034=0.034 [95%CI: 0.0000.07-0.000-0.07], $p=0.053$). This aligns with the interpretation that performance benefits emerge when systems achieve both redundancy (aligned toward a common goal) and synergistic integration (differentiated, complementary roles)—a form of functional and collective complexity (Varley et al., 2023; Luppi et al., 2024; Sterelny, 2007).
4.4 Experiments using Other Models
To assess generalizability of our findings, we repeat experiments with four other models: Llama 3.1 8B and 70B, Gemini 2.0 flash, and the reasoning model Qwen3 (see Appendix A.13 for full details). We find that the capacity for emergent synergy is robust across high-capability models: Llama 70B, Gemini, and Qwen3 achieved success rates on par with GPT 4.1 and exhibited strong evidence of emergence using both the emergence capacity and practical criteria. Consistent with our main results, these high-capability models achieved higher success rates, more differentiation, and stronger evidence of emergence in the Persona and ToM conditions compared to the Plain baseline. Conversely, the smaller Llama 8B largely failed to break oscillatory cycles and develop goal-directed cross-agent synergy due to its lower ToM reasoning reasoning capacity (Xiao et al., 2025). Crucially, our analysis of the reasoning model Qwen3 uncovered a distinct structural challenge we term paralysis under coordination ambiguity. Qwen3 agents enter infinite chain-of-thought loops when attempting to reconcile local binary search strategies with noisy group feedback. We provide a detailed taxonomy of these reasoning failure modes—including mutual mental modeling traps—in Appendix A.13, highlighting a critical frontier for research of reasoning models in multi-agent systems (Piedrahita et al., 2025).
@@ 5 Related Work
Shortly after the release of ChatGPT, the community began exploring multi-agent LLM systems, where multiple LLM agents interact with each other. One early example is Park et al. (2023), who simulated a small population of “individuals” in a The Sims-style environment with friends, houses, and jobs, showing the emergence of rich social behaviors such as coordinating a Valentine’s Day party. This was followed by multi-agent systems for complex tasks such as software development (Chen et al., 2023; Hong et al., 2024; Qian et al., 2023), healthcare (Li et al., 2024a), and other domains (Subramaniam et al., 2025; Ashery et al., 2025). Federated systems or “society of models” also gained attention (Juneja et al., 2024; Li et al., 2024b). These systems often achieve significant performance gains over single-agent baselines (Wu et al., 2023; Chen et al., 2023; Li et al., 2024b; Tao et al., 2024).
A recurring explanation for such gains invokes “greater-than-the-sum-of-its-parts” effects (Chen et al., 2023, p.1), often attributed to division of labor among differentiated agents (Subramaniam et al., 2025). Yet while this work is “inspired by human group dynamics” (Chen et al., 2023, p.1), it rarely evaluates whether the same principles that underlie effective (and ineffective) human groups also emerge in multi-agent systems. Two gaps stand out. First, human groups often coordinate by developing role specialization (Goldstone et al., 2024). Multi-agent systems often intuitively induce role specialization (such as “programmer”, “tester”, and “CEO”; Qian et al., 2023) yet have not systematically evaluated the impact on emergent shared cognition beyond win-rate comparisons. Second, groups only outperform individuals if their members specialize in contributing different pieces of information to the group task (Theiner, 2013; DeChurch and Mesmer-Magnus, 2010; Fazelpour and De-Arteaga, 2022; Nickerson and Zenger, 2004). Effective groups balance complementarity—members contributing distinct information—with redundancy, the alignment on shared goals (Theiner, 2013; DeChurch and Mesmer-Magnus, 2010; Fazelpour and De-Arteaga, 2022; Luppi et al., 2024; Tollefsen et al., 2013). Too much of either undermines performance. Unlike win-rate–centric evaluations—and tests of whether agents in multi-agent systems can sustain cooperation per se (Piedrahita et al., 2025; Piatti et al., 2024)—we propose falsification tests and null modeling, thus extending multi-agent frameworks with formal theory. While these principles likely apply to LLM-based collectives given the universality of the integration-segregation tradeoff (Tononi et al., 1994), it remains unclear whether agents develop differentiated roles, whether such roles complement each other, how to steer it with prompts, and what role ToM capacity of models plays for collaboration (Westby and Riedl, 2023; Kleiman-Weiner et al., 2025; Riedl and Weidmann, 2025).
@@ 6 Conclusion
Collectives of LLM agents are emerging as a powerful new paradigm, capable of tackling tasks that exceed the reach of any single model. Yet raw capability alone may not determine their effectiveness. Effective multi-agent systems depend on acting as integrated, cohesive units as opposed to loose aggregations of individual agents. In collective intelligence, this distinction often separates groups that merely average their parts from those that achieve true synergy (Riedl et al., 2021). This paper asks whether such coordination dynamics also matter for LLM collectives, what functional benefits they have, and whether we can design prompts, roles, or reasoning structures that foster such integrated yet synergistic behavior. The framework we developed connects human group cognition theory to LLM multi-agent systems, offering a novel conceptual bridge. Methodologically, we demonstrate how to operationalize group-level synergy in AI collectives using quantitative information-theoretic measures. We also provide specific design principles for creating productive LLM collectives that are aligned on “shared goals,” which can inform multi-agent orchestration tools, cooperative AI, and mixed human–AI teamwork. We further ground our information-theoretic framework in dynamical systems and game theory, showing that theory-of-mind interventions act as a control parameter that shifts multi-agent interactions from disordered to stable regimes with goal alignment. This work extends ToM research beyond standard false-belief tests to realistic collaborative tasks (Shapira et al., 2024; 2023) and points to ToM-related failure modes in reasoning models. Those insights will help us understand when multiple LLMs working together will be beneficial, why that is, and inform their design.
We note that evidence of higher-order synergy should not be interpreted as implying sophisticated cognition or consciousness. As conceptualized here, synergy is a structural property of part-whole relationships within multi-agent interactions. Such higher-order structures are known to arise in simple systems, including agents governed by reinforcement learning (Fulker et al., 2024) or via basic nonlinearities in contagion processes (Iacopini et al., 2019; Lee et al., 2025). Without attributing human-like cognition to the agents, the patterns of interaction we observe mirror well-established principles of collective intelligence in human groups (Riedl et al., 2021; DeChurch and Mesmer-Magnus, 2010): effective multi-agent performance requires both alignment on shared objectives and complementary contributions across members.
@@
This work focuses on developing a framework, characterizing emergence, and localizing it. While we explore the degree of goal alignment via a performance-related macro signal and performance effects of “early synergy”, the analysis linking synergy and redundancy to performance is challenging because synergy and redundancy are often co-dependent and emerge alongside performance. Future work should aim to more directly connect measures of synergy with performance. Despite a significant number of robustness tests, estimating entropy and establishing synergy is difficult. Furthermore, this work established results only using a single task—albeit one that is uniquely suited to studying synergy given how closely it mirrors ideal case of complementary behavior in a minimalist setting. More work will be necessary to convincingly establish when and how multi-agent LLM systems develop emergent synergy across tasks and measures. Given the small-data setting some of our information theoretic measures are limited to computing dynamic emergence on order $k=2$ which is bound to miss synergy of a higher order. Finally, this work points to the importance of ToM. LLMs capacity to act in a ToM-like manner appears crucial to achieve functional alignment in multi-agent systems. This should provide additional motivation for research on ToM in LLMs.